OpenStudy (anonymous):

An electron is projected horizontally at a speed of 1.21E+7m/s into an electric field. The field exerts a constant, vertically oriented force of 4.22E-16N on the particle. Determine the vertical distance by which the electron is deflected in the amount of time needed for it to have moved 28.0mm horizontally.

OpenStudy (anonymous):

You should find time t by using x=v*t Then, find the vertical acceleration by using F=m*a You can get the vertical distance by putting t and a into y=(a*t^2)/2